# A Quantum Perspective on Uniqueness Pairs for the Fourier Transform

### v1

@@ Abstract.
Kulikov, Nazarov, and Sodin recently introduced the notion of criticality to analyze when discrete subsets of the real line form uniqueness pairs for the Fourier transform, relying crucially on estimates derived from the Wirtinger–Poincaré inequality. In this work, drawing analogies from quantum mechanics, we propose an alternative approach to studying uniqueness sets in a Fourier-symmetric Sobolev space. Our method is based on eigenvalue estimates and offers a simplified perspective on the problem.
The Fourier-symmetric Sobolev space
$\mathcal{H}=\left\{f\>\Big|\>\int_{\mathbb{R}}(1+x^{2})|f(x)|^{2}\textnormal{d}x+\int_{\mathbb{R}}(1+k^{2})|\hat{f}(k)|^{2}\textnormal{d}k<\infty\right\}$
has recently attracted attention in the study of crystalline measures and sampling theory. In this note, we examine this space from a quantum-mechanical perspective, recovering recent results about uniqueness sets associated with the Fourier transform. We employ the “unitary convention” for the Fourier transform
$\hat{f}(k)=\mathcal{F}[f](k)=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}f(x)e^{-ikx}\textnormal{d}x,\quad f(x)=\mathcal{F}^{-1}[\hat{f}](x)=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}\hat{f}(k)e^{ikx}\textnormal{d}k.$
@@
It will be convenient to view $\mathcal{H}$ as a subspace of $L^{2}(\mathbb{R})$, by introducing the norm
@@
$H$ is the Hamiltonian operator of the classical quantum harmonic oscillator [S]. The eigenvalue-eigenvector equation $Hf=Ef$ is the Schrödinger equation of the system [S]. Our computation assumes that the second derivative of $f$ exists. However, for the construction of $\mathcal{H}$ this is immaterial.
@@
    $\mathcal{H}\subset L^{2}(\mathbb{R})$ and $\mathcal{H}$ is the domain and image of the self-adjoint and positive operator $H=\frac{1}{2}\left(x^{2}-\frac{\partial^{2}}{\partial x^{2}}\right)$.
@@
    $\phi_{n}(x)=\frac{1}{\sqrt{n!}}\left(\frac{1}{\sqrt{2}}(x-\partial_{x})\right)^{n}\phi_{0},\ n\geq 0,$
@@
    $H$ is self-dual under the Fourier transform: $\hat{H}=\mathcal{F}H\mathcal{F}^{-1}=H$.
@@
Similarly to the unconfined harmonic oscillator, the solutions of the confined system (0.1) have been studied exhaustively, e.g. [BS]. With the boundary conditions $H_{[a,b]}$ is a self-adjoint and positive definite operator on $L^{2}{[a,b]}$. We will write $E_{[a,b]}^{(n)}$ for the eigenvalues of the confined system and $\psi_{[a,b]}^{(n)}$ for the respective eigenvectors. It holds automatically that $E_{[a,b]}^{(n)}>E_{n}$ because the introduction of confinement can only lead to an increase of energy levels.
In the context of uniqueness sets of the Fourier transform and crystalline measures the concept of criticality has recently emerged as a means to characterize the supports of these measures. The following definition orients itself at [KNS]. Item (1) is taken from there [^1] The notions of super- and subcriticality studied in [KNS] are more general in that we restrict ourselves to the symmetric situation $(p,q)=(2,2)$, cf. Def. 2 in [KNS]. This is to achieve a lean presentation., whereas (2) is original to this note.
@@
Let $(\Lambda,M)$ be a pair of countable subsets of $\mathbb{R}$. Enumerate the elements of $\Lambda$ and $M$ in increasing order.
@@
An essential part in the proof of the theorem is played the Wirtinger–Poincaré inequality. We will use this theorem as a cornerstone to guide our further investigation. But rather than making use of the Wirtinger–Poincaré inequality, we will employ spectral methods to arrive at our estimates. Our definition of uniformly supercritical ground state energy is motivated by the following spectral estimate on the Rayleigh quotient of the confined Hamiltonian. It holds that
@@
Let the pair $(\Lambda,M)$ have uniformly supercritical ground energy. Then the only function $f\in\mathcal{H}$ such that $f(\lambda_{i})=\hat{f}(\mu_{i})=0$ for all $i\in\mathbb{Z}$ is $f=0$.
@@
We see three advantages in Theorem 5 over Theorem 4: 1) Obtaining lower estimates on the energy levels $E_{[a,b]}^{(n)}$ is straight forward. Their analytic form can be derived using perturbation theory. 2) Using eigenvalue bounds, the proof of Theorem 5 provides a direct route to sharpness, with the lower bound on the Rayleigh quotient attained by the ground-state eigenfunction. 3) The derivation of Theorem 5 is simpler than the one of Theorem 4 presented in [KNS].
We begin by studying the energy levels. Assuming $0\notin[a,b]$, the energy levels of the confined harmonic oscillators can be bounded by the respective energy levels of a particle in constant elevated box potential
@@
$\psi_{[a,b]}^{(n)}=\sqrt{\frac{2}{b-a}}\sin\left(\frac{\pi(n+1)(x-a)}{b-a}\right),$
@@
For the ground state energy of the confined Harmonic oscillators this implies
@@
    $\lim\inf\frac{\sqrt{E_{[\mu_{j},\mu_{j+1}]}^{(0)}}}{\max\{|\mu_{j}|,|\lambda_{j+1}|\}}$ $=\frac{\pi}{\lim\sup(\mu_{j+1}-\mu_{j})\max\{|\mu_{j}|,|\mu_{j+1}|\}}.$

### v2

# Spectral Criteria for Unique Signal Recovery from Two-Sided Sampling

@@ Abstract.
The identification of sampling sets that enable unique signal recovery is fundamental to many applications in signal processing and remains a central problem in mathematical analysis. Recent studies in the mathematical literature, particularly in the context of the Fourier transform and crystalline measures, have developed a theory that empowers signal recovery from two-sided sampling in both time and frequency domains. Kulikov, Nazarov, and Sodin introduced a method for identifying pairs of sets that enable unique recovery, based on functional inequalities of the Wirtinger-Poincaré type. In this work, we propose an alternative, spectral approach based on analogies with quantum mechanics. By relating uniqueness pairs to eigenvalue estimates of associated self-adjoint operators, our method offers a conceptually simpler and more flexible framework for studying signal recovery from two-sided sampling. Our approach extends naturally to other unitary transforms commonly used in signal processing. We demonstrate its effectiveness in the contexts of the Fractional Fourier transform and the Hankel transform.
The ability to reconstruct a function from its values on a discrete set lies at the heart of sampling theory, a cornerstone of both digital signal processing and applied mathematics. A fundamental application of this theory is the conversion between digital and analog signals. A central role in sampling theory is played by the concept of a uniqueness set, which generalizes the classical Nyquist rate of Whittaker-Kotelnikov-Shannon sampling theory beyond Paley-Wiener spaces. The characterization of uniqueness sets is closely tied to the geometry of the sampling space and the density of sampling points [S], making their study foundational to both theory and applications.
Recently, in the mathematical literature, particularly within the context of the Fourier transform and crystalline measures, the concept of a uniqueness pair has emerged. This pair consists of two sets: one in the real domain and the other in the Fourier domain, with the property that if values are specified for an interpolating function at these sets, this function is determined uniquely, see [RV], and also [RS, KNS]. Beyond its mathematical significance, this concept has practical implications in engineering, as it enables the unique recovery of a signal from two-sided sampling - i.e., when samples are concurrently taken in both the real and Fourier domains. Achieving this typically requires that the sampling densities in both domains surpass certain critical thresholds, the exact form of which depends on the geometry of the underlying spaces.
Current approaches to determining these recovery thresholds rely crucially on estimates derived from the Wirtinger–Poincaré inequality [KNS]. Motivated by analogies with quantum mechanics, we propose an alternative spectral approach to the analysis of uniqueness pairs. Our framework offers a natural and flexible method for addressing uniqueness problems for the Fourier transform and extends directly to a broad class of unitary transformations on the real line.
Beyond the classical Fourier transform, several related unitary transforms play a fundamental role in signal processing and engineering applications. We extend the idea of two-sided sampling to these settings and we provide a methodology to study uniqueness pairs for such transformations. To illustrate the effectiveness of our approach, we consider the Fractional Fourier and the Hankel transforms. The Fractional Fourier transform provides a continuous interpolation between the time and frequency domains and is widely used in optical signal analysis, radar imaging, and time-frequency filtering [OZK]. The Hankel transform arises naturally in problems with radial symmetry, including wave propagation, diffraction, and medical imaging such as computer tomography [KS].
1. Uniqueness pairs for the Fourier transform
We begin by considering the Fourier-symmetric Sobolev space
$\mathcal{H}=\left\{f\>\Big|\>\int_{\mathbb{R}}(1+x^{2})|f(x)|^{2}\textnormal{d}x+\int_{\mathbb{R}}(1+k^{2})|\hat{f}(k)|^{2}\textnormal{d}k<\infty\right\},$
which arises naturally in the theory of modulation spaces - function spaces used in time-frequency analysis to quantify the joint localization of a signal in both time and frequency; see, e.g., [EG]. More recently, $\mathcal{H}$ has also attracted attention in the study of crystalline measures and two-sided sampling theory [KNS, Z]. In this section, we examine $\mathcal{H}$ from a quantum-mechanical perspective and recover several recent results concerning uniqueness pairs associated with the Fourier transform. We adopt the unitary convention for the Fourier transform, as it provides conceptual clarity and facilitates the generalization of the notion of uniqueness pairs to other unitary transformations:
$\hat{f}(k)={F}[f](k)=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}f(x)e^{-ikx}\textnormal{d}x,\quad f(x)={F}^{-1}[\hat{f}](x)=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}\hat{f}(k)e^{ikx}\textnormal{d}k.$
@@ 1. Uniqueness pairs for the Fourier transform
We note that the convention used in [KNS] is not unitary. It will be convenient to view $\mathcal{H}$ as a subspace of $L^{2}(\mathbb{R})$, by introducing the norm
@@
$H$ is the Hamiltonian operator of the classical quantum harmonic oscillator [NS]. The eigenvalue-eigenvector equation $Hf=Ef$ is the Schrödinger equation of the system [NS]. Our computation assumes that the second derivative of $f$ exists. However, for the construction of $\mathcal{H}$ this is immaterial.
@@
    $\mathcal{H}\subset L^{2}(\mathbb{R})$ and $\mathcal{H}$ is the domain and image of an essentially self-adjoint and positive operator $H=\frac{1}{2}\left(x^{2}-\frac{\partial^{2}}{\partial x^{2}}\right)$.
@@
    $\phi_{n}(x)=\frac{1}{\sqrt{n!}}\left(\frac{1}{\sqrt{2}}(x-\partial_{x})\right)^{n}\phi_{0}(x),\ n\geq 0,$
@@
    $H$ is self-dual under the Fourier transform: $\hat{H}={F}H{F}^{-1}=H$.
@@
Similarly to the unconfined harmonic oscillator, the solutions of the confined system (1.1) have been studied exhaustively, e.g. [BS]. With the boundary conditions $H_{[a,b]}$ is a self-adjoint and positive definite operator on $L^{2}{[a,b]}$. We will write $E_{[a,b]}^{(n)}$ for the eigenvalues of the confined system and $\psi_{[a,b]}^{(n)}$ for the respective eigenvectors. It holds automatically that $E_{[a,b]}^{(n)}>E_{n}$ because the introduction of confinement can only lead to an increase of energy levels. The following definition follows [KNS]. Item (1) is taken from there [^1] The notions of super- and subcriticality studied in [KNS] are more general in that we restrict ourselves to the symmetric situation $(p,q)=(2,2)$, cf. Def. 2 in [KNS]. This is to achieve a lean presentation., whereas (2) is original to this note.
@@
Let $(\Lambda,M)$ be a pair of countable subsets of $\mathbb{R}$. Enumerate the elements of $\Lambda$ and $M$ in increasing order, denoting them by $(\lambda_{j})_{j\in\mathbb{Z}}$ and $(\mu_{j})_{j\in\mathbb{Z}}$, respectively.
@@
The elegant proof presented in [KNS] relies crucially on the Wirtinger–Poincaré inequality. A closely related argument, involving the same inequality, has appeared in the context of irregular sampling in [G]. We will use this theorem as a cornerstone to guide our further investigation. But rather than making use of the Wirtinger–Poincaré inequality, we will employ spectral methods to arrive at our estimates. Our definition of uniformly supercritical ground state energy is motivated by the following spectral estimate on the Rayleigh quotient of the confined Hamiltonian. It holds that
@@
Let the pair $(\Lambda,M)$ have uniformly supercritical ground state energy. Then the only function $f\in\mathcal{H}$ such that $f(\lambda_{i})=\hat{f}(\mu_{i})=0$ for all $i\in\mathbb{Z}$ is $f=0$.
@@
We see three advantages in Theorem 5 over Theorem 4: 1) Lower bounds on the energy levels $E_{[a,b]}^{(n)}$ are straightforward to obtain. Their analytic form can be derived using perturbation theory. 2) The proof of Theorem 5, via eigenvalue bounds, gives a direct route to establishing sharpness, with the lower bound on the Rayleigh quotient attained by the ground-state eigenfunction. 3) The derivation of Theorem 5 is simpler than that of Theorem 4 as presented in [KNS], and it naturally extends to unitary transformations beyond the Fourier transform.
We begin by studying the energy levels. Assuming $0\notin[a,b]$, the energy levels of the confined harmonic oscillator can be bounded by the respective energy levels of a particle in constant elevated box potential
@@
$\psi_{[a,b]}^{(n)}=\sqrt{\frac{2}{b-a}}\exp\left(i\frac{\pi(n+1)(x-a)}{b-a}\right),$
@@
For the ground state energy of the confined harmonic oscillator this implies
@@
    The notions of supercriticality and supercriticality of the ground state energy are equivalent
@@
    $\lim\inf\frac{\sqrt{E_{[\mu_{j},\mu_{j+1}]}^{(0)}}}{\max\{|\mu_{j}|,|\mu_{j+1}|\}}$ $=\frac{\pi}{\lim\sup(\mu_{j+1}-\mu_{j})\max\{|\mu_{j}|,|\mu_{j+1}|\}}.$
    The analogous statement holds also for subcriticality and supercriticality of the ground state energy.